OpenStudy (anonymous):

hello everyone! the name is Mikel Formanes, a BS-Mathematics student from SLSU in Lucban, Philippines. Can someone help me about a problem regarding Partial Derivatives...

OpenStudy (anonymous):

whats a problem?

hartnn (hartnn):

@maikeL Hi, \(\Huge \mathcal{\text{Welcome To OpenStudy} \ddot\smile} \) can you post a particular problem you have doubt with ? we'll try our best to help you out.

OpenStudy (shubhamsrg):

http://gyazo.com/babac224247a886dfcc9933b92349a81 @Ayeshaafzal ^_^

OpenStudy (anonymous):

graphing a function of one variable is quite easy, but i had a hard time analyzing and graphing function of two variables..

OpenStudy (anonymous):

Definition: a function of two variables, x and y, assigns a unique real number ƒ(x,y) to each point (x,y) in the xy plane. Q: how will i graph the fcn. ƒ(x,y)= √(1-x²-y²) ?

hartnn (hartnn):

ƒ(x,y)= √(1-x²-y²) = z so, x^2+y^2 +z^2 = 1 from this , you can figure out that its a 3-dimensional plot. its a sphere of radius = 1 and only part above xy plane (because its +√(1-x²-y²) and not -√(1-x²-y²) )

OpenStudy (unklerhaukus):

when y=0 f=√(1-x^2) when x=0 f=√(1-y^2) can you plot these intersection of the function ?

hartnn (hartnn):

standard equation of sphere, x^2+y^2+z^2 = r^2 radius = r

OpenStudy (anonymous):

@hartnn so it's a sphere centered at the origin with radius 1 right.. z must be greater than or equal to 0 so its a sphere with just the upper hemisphere.

hartnn (hartnn):

yes, thats what it is.

OpenStudy (unklerhaukus):

\[\huge\color{red}\checkmark{}\]

OpenStudy (anonymous):

just for the record, thanks for your responses @hartnn and @UnkleRhaukus . :)

hartnn (hartnn):

welcome ^_^ since you are new here, if you need any assistance browsing this site , you can ask me :)

OpenStudy (anonymous):

are you familiar with the term level curves?

OpenStudy (unklerhaukus):